# Constraining the dispersion measure redshift relation with simulation-based inference

### v1

@@ 4 Inference
Table 1: Parameters fitted in the inference with prior ranges and fiducial values for the data compression.
@@ 4.1 Simulation-based inference
Approximate Bayesian Calculation (ABC) addresses this issue by comparing data from forward simulation with observation with some arbitrary margin (Rubin, 1984; Pritchard et al., 1999). Only the accepted simulations are considered for posterior inference, which leads to critical drawbacks such as a high rejection rate with decreasing margin and the curse of dimensionality (Sisson et al., 2018). Improvements upon this basic rejection sampling include MCMC-ABC (Marjoram et al., 2003) and Sequential Monte Carlo ABC (SMC-ABC) (Bonassi & West, 2015), which only partially solve the rejection rate and do not address the high dimensionality problem.
@@
The fundamental idea behind all of these methods is to approximate the posterior, $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$, with a conditional density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$, where $\phi$ are the trainable parameters. Typically, the parameters, $\boldsymbol{\theta}$, are sampled from the full prior range, which is inefficient. There exist two schools of SBI. Amortised methods yield a posterior that can be utilised for various observations without the need for retraining, whereas sequential methods concentrate the inference on a specific observation to enhance simulation efficiency. Here we focus on the sequential-(S)NPE methods refer to using a proposal as a subset of the prior, $\tilde{p}\boldsymbol{\theta})\subseteq p(\boldsymbol{\theta})$. Initially, the proposal is equal to the prior and in subsequent rounds we update the proposal with the approximate posterior, improving the efficiency of the inference process. The predicament is that this procedure does not recover the true posterior, but rather an approximate posterior:
@@
The problem to solve is the extra ratio, $\tilde{p}(\boldsymbol{\theta})/p(\boldsymbol{\theta})$. The way each algorithm addresses this issue is what differentiates the various SNPE methods. For instance, SNPE-A maintains a closed-form solution by restricting the choice of the prior, proposal and the density estimator to either a uniform or Gaussian distribution. The training is carried out by minimising the loss function or the negative log-likelihood of the joint probability of the data and parameter defined via
@@
Other methods like SNPE-B offer a solution by including the ratio inside the loss function, whereas SNPE-C incorporates the ratio into the density estimator and finally, APT normalises the density estimator by using uniform subsets of the prior called atoms. This development allows any arbitrary distribution choice for the prior, proposal and the density estimator. For an extensive technical review of the shortcomings and the subsequent developments of the various SNPE methods, the reader is referred to the following articles and the references therein (Durkan et al., 2020; Lueckmann et al., 2021; Xiong et al., 2023).
The latest development, TSNPE, builds upon the APT method and truncates the prior based on the $1-\epsilon$ mass ($\epsilon$ being an arbitrary margin) of the highest probability region (HPR) of the approximate posterior. In other words, the subsequent rounds only consider the prior range which is most probable to produce the posterior. This prescription maintains the flexibility of the distribution choice while also accounting for the posterior mass outside the prior boundaries, an issue of the APT method (Deistler et al., 2022).
@@
The various SNPE methods tackle posterior estimation by either changing the loss function or restricting the choice of distribution. Without a general solution, the choice of algorithm affects the performance and accuracy of the analysis. Our tests reveal that TSNPE, due to the truncation, is faster and allows for a more flexible distribution choice for prior and density estimators as compared to SNPE-A. As the field of NN is projected to experience rapid growth, the choice of the best algorithm is expected to change as well; refer to (Cranmer et al., 2020) for a detailed review of the status of SBI at the beginning of this decade.
@@ 4.2 Implementation
The simulated data in our case are the compressed DM values of FRBs as described in Section 3.2. The priors are defined in Table 1. For the analysis, we use 12 FRBs with host identification as given in Footnote 4, where amongst others we list the observed DMs and their references.
@@
In each round of TSNPE, the density estimator learns the mapping between the simulated data and the corresponding parameters. The observed DM is introduced post training, which helps evaluate an approximate posterior for that round. This approximate posterior is then utilised to calculate the HPR region and the prior range for the next round is truncated based on that. As the number of rounds is increased, the approximate posterior converges to the true posterior. Introducing the observed DM is necessary to propel the truncation in the direction of the true posterior. With parameters, simulated data and observation, these are the three components required for the SBI analysis. We implement this using LtU-ILI package for a quick setup with different neural embedding (Ho et al., 2024). We have used the TSNPE algorithm for a total of 1500 simulations in 10 rounds. The initial simulations contain the full prior range and the subsequent truncation is only influenced by the data. As LtU-ILI does not natively support TSNPE yet, we run TSNPE, save the data and then use it. The choice of the number of simulations in each round is dictated by the number of epochs the NN takes for training. We limit the number of epochs to 50 to avoid overfitting and 150 simulations in each round provide ample data for the density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$ to approximate the posterior by adjusting the trainable parameter $\phi$.
As for the NN architecture, we use two different types; mixture density network (MDN) and masked autoregressive flows (MAF). These architectures aim to produce an arbitrary yet tractable PDF with trainable parameters to be used as the density estimator. The complexity of data mandates arbitrariness, while the backpropagation in the training process requires tractability. MDN, as the name suggests, is the weighted sum of many Gaussian distributions defined as (Bishop, 1994)
@@
For the MDN embedding, we use 50 hidden layers with 10 components in each layer, while the MAF net is constructed using 20 hidden layers with 8 transformations in each layer. When the training is complete, we acquire a mapping (effectively a KDE) from the joint space of parameters to the distribution $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$ without the usage of an explicit likelihood function. This posterior can then be sampled by standard MCMC techniques.
@@ 5.1 Coverage Test
To check the consistency of the posterior distributions, we use the multivariate coverage test called TARP. Coverage probability measures how frequently the estimated posterior contains the true parameter value, and can be used to check the consistency of the posterior (Guo et al., 2017; Hermans et al., 2021). Once we have obtained the posterior, TARP measures the expected coverage probability of random posterior samples within a given credibility level of the learned posterior empirically. In more technical terms, for any instance of the prior $\boldsymbol{\theta}^{\star}$ and the corresponding simulated data $\boldsymbol{\mathrm{d}}^{\star}$, samples are drawn from the posterior as $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}^{\star}})$. Then, a circle is drawn centred at a random reference point $\boldsymbol{\theta}_{\mathrm{r}}$ with radius $|\boldsymbol{\theta}^{\star}-\boldsymbol{\theta}_{\mathrm{r}}|$ and the fraction of posterior samples inside the circle is calculated. A higher fraction implies the posterior is more accurate. The fraction is calculated multiple times to get a statistical estimate of the coverage.
This random sampling is necessary to validate the static HPR region in TSNPE, which is prone to having blind spots. We use the learnt posterior to draw 1000 samples at random parameter points for the TARP test implemented within the LtU-ILI. We also bootstrap the test 100 times at each credibility level, considering the process is random by definition.
Figure 3: TARP coverage for the fiducial case of log-normal field and log-normal host is shown with $N_{\mathrm{side}}=4096$. Credibility level refers to the fraction of the total PDF of the final posterior in Figure 2 being considered, which intuitively goes from 0 to 1. The expected coverage is the fraction of the posterior samples with a lower posterior probability than the best estimate for that corresponding credibility. The diagonal line is the ideal relation, as increasing the credibility level linearly increases the expected coverage. For example, at a 50% credibility level, we expect at least 50% of the samples to have coverage; anything below 50% implies the samples are not well covered, i.e. the posterior is biased and similarly a value greater than 50% implies that the posterior is conservative. The dark and light-shaded regions are the one and two-sigma error bars from 100 bootstrappings. Our model, the solid blue line, is accurate as the ideal line is within the $1\sigma$ error bar.
@@ 5.2 Goodness-of-Fit
where E denotes the expectation value and Cov is the covariance at the best-fit cosmology. $\boldsymbol{t}_{i}$ is the score-compressed summary statistic of the $i$-th realisation. Specifically, we create 1000 new forward simulations at the best-fit values to calculate the mean data vector and the covariance. The $\chi^{2}$ values from the simulations are plotted as a histogram and compared with the $\chi^{2}$ value for the observed DM. The accuracy of the model is assessed through the probability mass of the simulated $\chi^{2}$ that lies beyond this observed $\chi^{2}$, which we call the probability to exceed (PTE).
@@ 6 Results
In this section, we present the major result of our analysis. We consider three parameters for the inference. The first one is the amplitude of the DM-$z$ relation in Equation 7 as a scale. This encapsulates the $\Lambda$CDM parameters, for which we use the Planck cosmological results from (Planck Collaboration et al., 2020), i.e. $\mathcal{A}=1$ refers to the input cosmology and this $\mathcal{A}$ is varied in different simulations. The motivation behind only varying the amplitude of the DM-redshift relation is that the number of FRBs currently available is not able to discriminate between different expansion histories. The remaining two parameters are the median and the scale of the log-normal host distribution, which we denote as $\mathrm{DM}_{\mathrm{host}}$ and $\sigma_{\mathrm{LN}}$.
As for the model, there are two components in our simulations that vary. First, the $\mathrm{DM}_{\mathrm{LSS}}$ part is modelled using either a log-normal or a Gaussian realisation of the underlying electron density field via GLASS. These are denoted by ‘Field: log-normal’ and ‘Field: Gaussian’ respectively. Similarly, we have three host distributions; log-normal, truncated Gaussian and Gamma. With this setup, we have six possible combinations of field and host. That being said, we use a log-normal field and a log-normal host distribution as our fiducial setting. For this, we will discuss and validate the results in detail and use the high resolution setting of $N_{\mathrm{side}}=4096$. For the other combinations, we use a lower resolution $N_{\mathrm{side}}=512$ in order to limit computation time. In the following, we start with the fiducial model and then observe how the changes in field and host affect the results.
@@ 6.1 Fiducial model and validation
For the fiducial model, we assume the electron density field in GLASS to be log-normal and sample the host DM from a log-normal distribution. The decision is primarily motivated by results from simulations (see e.g. Zhang et al., 2021; Theis et al., 2024). In the subsequent sections, we present the comparisons of the six different models, which support our decision.
With the simulator set to the desired configuration and priors defined in Table 1, the SBI pipeline is run with the TSNPE algorithm, which returns the posterior distribution for the free parameters. In Figure 2, we show the marginalised contour plots using ChainConsumer (Hinton, 2016). The forward simulations have the resolution parameter $N_{\mathrm{side}}=4096$. As mentioned at the beginning of this section, there are three parameters that we constrain; $\mathcal{A}$ as a scale for the input cosmology, $\mathrm{DM}_{\mathrm{host}}$ as the median and $\sigma_{\mathrm{LN}}$ as the scale of the log-normal host distribution. The value for the scale parameter $\mathcal{A}$ is $0.89^{+0.31}_{-0.29}$ at $68\%$ confidence, which is consistent with unity or the input fiducial Planck cosmology. The median and the scale of the log-normal host along with their $68\%$ confidence intervals are given as $\mathrm{DM}_{\mathrm{host}}=223^{+108}_{-95}\;\mathrm{pc\;cm}^{-3}$ and $\sigma_{\mathrm{LN}}=0.76^{+0.37}_{-0.44}$. This $\sigma_{\mathrm{LN}}$ is further converted into the standard deviation from the definition of the variance in Equation 19. Then, we write $\sigma=263_{-113}^{+91}$. As expected, we find a strong anti-correlation between $\mathcal{A}$ and $\mathrm{DM}_{\mathrm{host}}$ since they both enter the observed DM in an additive fashion. The width of the host contribution is not degenerate with the other parameters, as it rather reduces the error on each measurement than changing the signal.
Our constraints of $\sigma_{\mathrm{LN}}$ seem to be slightly prior driven, as can be seen from the posterior hitting the prior boundaries of $\sigma_{\mathrm{LB}}$ in Figure 2. However, there is clearly some constraining power in the data on the shape of the host distribution. In general, we find excellent agreement with previous work (Macquart et al., 2020; James et al., 2022; Hagstotz et al., 2022b; Reischke & Hagstotz, 2023a; Khrykin et al., 2024b).
@@
Additionally, we check for the goodness of fit of the model and the result is shown in Figure 4. The histograms are the $\chi^{2}$ values from our simulations at best-fit values of the parameters, and the vertical line corresponds to the $\chi^{2}$ evaluated at the observed DM. The PTE for the fiducial model is 0.3 (corresponding to a $p$-value), which means that the model is a good fit. We can thus concur that the ‘log-normal field and log-normal host’ model is an excellent fit for the DM of FRBs.
@@
With the SBI framework, all the correlations are taken into account without defining any likelihood function. Nonetheless, we can still recover the likelihood from the simulated DMs for the best-fit values. We therefore run our pipeline 1000 times at the best-fit value and plot the corresponding distribution of DM. In Figure 5 we show the DM likelihoods for two FRBs with varying redshift, $z=0.1178$ and $z=0.66$. The first observation is the non-Gaussian nature of the distribution, with a tendency towards higher DMs. This behaviour emerges from the host and LSS DM contribution. If the electron density of the host halo where the FRB progenitor resides is high, the signal experiences higher dispersion. That seems to be especially the case for FRB 20190520B, with an observed $\mathrm{DM}=1202\;\mathrm{pc\;cm}^{-3}$ at a relatively lower redshift of $z=0.241$ as shown in Footnote 4. The high DM is due to the local contribution of the dwarf host galaxy, identified as J160204.31-111718.5 (Ocker et al., 2022; Niu et al., 2022; Yan et al., 2024). The implication is that a Gaussian likelihood assumption in the inference can introduce biases, as it does not consider the physical effect of variable travel distances and local environments of the host. Furthermore, we observe that the mean shifts towards higher DMs in keeping with the DM-$z$ relation. We also see that the width of the distribution seems to be getting slightly smaller. This might be counter-intuitive at first, as the LSS component should increase with redshift, hence causing a broader distribution. However, the host contribution is scaled down with $1/(1+z)$ in physical coordinates. At the same time, the density of the electron density field reduces as well with increasing redshift [^2] It is still an additive effect, of course.. Thus, those two counter-acting effects can lead to a decreasing width of the likelihood as a function of redshift. The strength of this effect itself depends on the parameters of the model, in particular on the parameters of the host contribution. To understand this a bit better, let us consider a simple toy model. We call the variance of the host at redshift zero $\sigma^{2}_{\mathrm{host,0}}$ and label the variance of the LSS at $z=1$ as $\sigma^{2}_{\mathrm{LSS,1}}$. For linear structure growth in a matter-dominated Universe, one has $\delta_{\mathrm{e}}\propto a$. The LSS variance scales roughly linear with redshift in this case (Reischke & Hagstotz, 2023b). Therefore, the total variance is given by
@@
Finally, we show the model prediction at the global maximum posterior together with the likelihood for all data points in Figure 6. Note that the model prediction includes all components of the DM in Equation 1. Therefore, the relation between DM and redshift is not necessarily monotonous. The medians of the simulated DMs are shown in cyan dots for the redshifts of the FRBs. The colour bars represent the different percentiles. They include the observed data (black cross) within the darker shades, i.e. 25 to 75 percentiles. We would like to remark that those percentiles are in principle correlated due to the LSS contribution. However, this correlation is not important for the number of FRBs considered here (Reischke & Hagstotz, 2023b). It is again apparent that the long tails of the likelihood are required to explain the large scatter in DM values. We furthermore investigated the response of the inference to leaving out FRB20190520B, which has a DM > 1000 but is located at a low redshift. Our findings show that the host contribution responds with a lower median and width by roughly 10 and 20 percent respectively. This of course is still fully consistent within the error bars. It shows, however, that this particular FRB increases the values of the inferred host contribution. Lastly, we recover the mean DM-$z$ relation as well by connecting the cyan dots.
@@
Figure 5: Likelihood of the DM for two FRBs at different redshifts as evaluated from the simulations. The blue line is for a lower redshift of $z=0.1178$, while the red line is for $z=0.66$. The non-Gaussian nature is obvious along with the distinct long tail which accounts for the high DM values originating from the disk of the host halo.
6.2 Variations in the LSS component

### v2

@@ 4 Inference
Table 1: Parameters fitted in the inference with prior ranges (flat) and fiducial values for the data compression.
@@ 4.1 Simulation-based inference
The fundamental idea behind all of these methods is to approximate the posterior, $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$, with a conditional density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$, where $\phi$ are the trainable parameters. Typically, the parameters, $\boldsymbol{\theta}$, are sampled from the full prior range, which is inefficient. There exist two schools of SBI. Amortised methods yield a posterior that can be utilised for various observations without the need for retraining, whereas sequential methods concentrate the inference on a specific observation to enhance simulation efficiency. Note that these two methods are not strictly mutually exclusive. Even though sequential SBI is tailored to a specific observation, the resulting posterior can be utilised in an amortised manner. If the new observation for the amortised case is not significantly different from the one used in sequential training, the older network can be applied with additional training.
Here, we focus on the sequential-(S)NPE methods, which refer to using a proposal as a subset of the prior, $\tilde{p}\boldsymbol{\theta})\subseteq p(\boldsymbol{\theta})$. Initially, the proposal is equal to the prior and in subsequent rounds, we update the proposal with the approximate posterior, improving the efficiency of the inference process. The predicament is that this procedure does not recover the true posterior, but rather an approximate posterior:
@@
The problem to solve is the extra ratio, $\tilde{p}(\boldsymbol{\theta})/p(\boldsymbol{\theta})$. The way each algorithm addresses this issue is what differentiates the various SNPE methods. For instance, SNPE-A maintains a closed-form solution by restricting the choice of the prior, proposal and density estimator to either a uniform or Gaussian distribution. The training is carried out by minimising the loss function or the negative log-likelihood of the joint probability of the data and parameter defined via
@@
Other methods like SNPE-B offer a solution by including the ratio inside the loss function, whereas SNPE-C incorporates the ratio into the density estimator and finally, APT normalises the density estimator by using uniform subsets of the prior called atoms. This development allows any arbitrary distribution choice for the prior, proposal and density estimator. For an extensive technical review of the shortcomings and the subsequent developments of the various SNPE methods, the reader is referred to the following articles and the references therein (Durkan et al., 2020; Lueckmann et al., 2021; Xiong et al., 2023).
The latest development, TSNPE, builds upon the APT method and truncates the prior based on the $1-\epsilon$ mass ($\epsilon$ being an arbitrary margin) of the highest probability region (HPR) of the approximate posterior. In other words, the subsequent rounds only consider the prior range, which is most probable to produce the posterior. This prescription maintains the flexibility of the distribution choice while also accounting for the posterior mass outside the prior boundaries, an issue of the APT method (Deistler et al., 2022).
@@
The various SNPE methods tackle posterior estimation by either changing the loss function or restricting the choice of distribution. Without a general solution, the choice of algorithm affects the performance and accuracy of the analysis. Our tests reveal that TSNPE, due to the truncation, is faster and allows for a more flexible distribution choice for prior and density estimators as compared to SNPE-A. As the field of NNs is projected to experience rapid growth, the choice of the best algorithm is expected to change as well; refer to (Cranmer et al., 2020) for a detailed review of the status of SBI at the beginning of this decade.
@@ 4.2 Implementation
The simulated data in our case are the compressed DM values of FRBs as described in Section 3.2. We use uniform priors and the ranges are defined in Table 1, where the limits are motivated by the results from Hagstotz et al. (2022b); Reischke et al. (2022). The priors are all wide enough, especially for $\mathcal{A}$ and $\mathrm{DM}_{\mathrm{host}}$, that the data have to constrain them. Since $\sigma_{\mathrm{LN}}$ controls the shape of the distribution, it will be more dominated by the prior as this would require more FRBs. However, increasing $\sigma_{\mathrm{LN}}$ above the provided prior bound would create a host PDF whose peak is at very low DM; this has not been observed in simulations (Theis et al., 2024; Medlock et al., 2024). For the analysis, we use 12 FRBs with host identification as given in Footnote 3, where, amongst others, we list the observed DMs and their references.
@@
In each round of TSNPE, the density estimator learns the mapping between the simulated data and the corresponding parameters. The observed DM is introduced post-training, which helps evaluate an approximate posterior for that round. This approximate posterior is then utilised to calculate the HPR region and the prior range for the next round is truncated based on that. As the number of rounds is increased, the approximate posterior converges to the true posterior. Introducing the observed DM is necessary to propel the truncation in the direction of the true posterior. With parameters, simulated data and observation, these are the three components required for the SBI analysis. We implement this using the LtU-ILI package for a quick setup with different neural embedding (Ho et al., 2024). We have used the TSNPE algorithm for a total of 1500 simulations in 10 rounds. The initial simulations contain the full prior range and the subsequent truncation is only influenced by the data. As LtU-ILI does not natively support TSNPE yet, we run TSNPE, save the data and then use it. The choice of the number of simulations in each round is dictated by the number of epochs the NN takes for training. We limit the number of epochs to 50 to avoid overfitting and 150 simulations in each round provide ample data for the density estimator $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$ to approximate the posterior by adjusting the trainable parameter $\phi$.
As for the NN architecture, we use two different types: mixture density network (MDN) and masked autoregressive flows (MAF). These architectures aim to produce an arbitrary yet tractable PDF with trainable parameters to be used as the density estimator. The complexity of data mandates arbitrariness, while the backpropagation in the training process requires tractability. MDN, as the name suggests, is the weighted sum of many Gaussian distributions defined as (Bishop, 1994)
@@
For the MDN embedding, we use 50 hidden layers with 10 components in each layer, while the MAF net is constructed using 20 hidden layers with 8 transformations in each layer. The network is trained to minimise the negative log-likelihood or the Kullback–Leibler (KL) divergence between the approximate and the true posterior. For our case, this minimisation is achieved with $\sim$ 800 simulations. When the training is complete, we acquire a mapping (effectively a KDE) from the joint space of parameters to the distribution $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}})$ without the usage of an explicit likelihood function. This posterior can then be sampled by standard MCMC techniques (here we use emcee introduced in Foreman-Mackey et al., 2013). We use 3 walkers and they all converge consistently.
@@ 5.1 Coverage Test
To check the consistency of the posterior distributions, we use the multivariate coverage test called TARP. Coverage probability measures how frequently the estimated posterior contains the true parameter value and can be used to check the consistency of the posterior (Guo et al., 2017; Hermans et al., 2021). Once we have obtained the posterior, TARP measures the expected coverage probability of random posterior samples within a given credibility level of the learned posterior empirically. In more technical terms, for any instance of the prior $\boldsymbol{\theta}^{\star}$ and the corresponding simulated data $\boldsymbol{\mathrm{d}}^{\star}$, samples are drawn from the posterior as $p(\boldsymbol{\theta}|\boldsymbol{\mathrm{d}^{\star}})$. Then, a circle is drawn centred at a random reference point $\boldsymbol{\theta}_{\mathrm{r}}$ with radius $|\boldsymbol{\theta}^{\star}-\boldsymbol{\theta}_{\mathrm{r}}|$ and the fraction of posterior samples inside the circle is calculated. A higher fraction implies the posterior is more accurate. The fraction is calculated multiple times to get a statistical estimate of the coverage.
This random sampling is necessary to validate the static HPR region in TSNPE, which is prone to having blind spots. We use the learnt posterior to draw 1000 samples at random parameter points for the TARP test implemented within the LtU-ILI. All three parameters in our analysis, $\mathcal{A},\,\mathrm{DM}_{\mathrm{host}}$ and $\sigma_{\mathrm{LN}}$, are kept free and their prior ranges are defined in Table 1. We also bootstrap the test 100 times at each credibility level, considering the process is random by definition. We elaborate on the result of the coverage test as shown in Figure 2 in Section 6.
Figure 2: TARP coverage for the fiducial case of log-normal field and log-normal host is shown with $N_{\mathrm{side}}=4096$. Credibility level refers to the fraction of the total PDF of the final posterior in Figure 3 being considered, which intuitively goes from 0 to 1. The expected coverage is the fraction of the posterior samples with a lower posterior probability than the best estimate for that corresponding credibility. The diagonal line is the ideal relation, as increasing the credibility level linearly increases the expected coverage. For example, at a 50% credibility level, we expect at least 50% of the samples to have coverage; anything below 50% implies the samples are not well covered, i.e. the posterior is biased and similarly, a value greater than 50% implies that the posterior is conservative. The dark and light-shaded regions are the one and two-sigma error bars from 100 bootstrappings. Our model, the solid blue line, is accurate as the ideal line is within the $1\sigma$ error bar.
@@ 5.2 Goodness-of-Fit
where E denotes the expectation value and Cov is the covariance at the best-fit cosmology. $\boldsymbol{t}_{i}$ is the score-compressed summary statistic of the $i$-th realisation. Specifically, we create 1000 new forward simulations at the best-fit values to calculate the mean data vector and the covariance. The $\chi^{2}$ values from the simulations are plotted as a histogram and compared with the $\chi^{2}$ value for the observed DM. The resulting figure is shown in Figure 4 and explained in Section 6. The accuracy of the model is assessed through the probability mass of the simulated $\chi^{2}$ that lies beyond this observed $\chi^{2}$, which we call the probability to exceed (PTE).
@@ 6 Results
In this section, we present the major results of our analysis. We consider three parameters for the inference. The first one is the amplitude of the DM-$z$ relation in Equation 7 as a scale. This encapsulates the $\Lambda$CDM parameters, for which we use the Planck cosmological results from (Planck Collaboration et al., 2020), i.e. $\mathcal{A}=1$ refers to the input cosmology and this $\mathcal{A}$ is varied in different simulations. The motivation behind only varying the amplitude of the DM-redshift relation is that the number of FRBs currently available is not able to discriminate between different expansion histories. In other words, we can only aim to infer $H_{0}$ but not $E(z)$ (see e.g. Macquart et al., 2020b; Hagstotz et al., 2022b; James et al., 2022b). The remaining two parameters are the median and the scale of the log-normal host distribution, which we denote as $\mathrm{DM}_{\mathrm{host}}$ and $\sigma_{\mathrm{LN}}$.
As for the model, there are two components in our simulations that vary. First, the $\mathrm{DM}_{\mathrm{LSS}}$ part is modelled using either a log-normal or a Gaussian realisation of the underlying electron density field via GLASS. These are denoted by ‘Field: log-normal’ and ‘Field: Gaussian’ respectively. Similarly, we have three host distributions: log-normal, truncated Gaussian and Gamma. With this setup, we have six possible combinations of field and host. That being said, we use a log-normal field and a log-normal host distribution as our fiducial setting. For this, we will discuss and validate the results in detail and use the high resolution setting of $N_{\mathrm{side}}=4096$. For the other combinations, we use a lower resolution $N_{\mathrm{side}}=512$ in order to limit computation time. In the following, we start with the fiducial model and then observe how the changes in field and host affect the results.
@@ 6.1 Fiducial model and validation
For the fiducial model, we assume the electron density field in GLASS to be log-normal and sample the host DM from a log-normal distribution. The decision is primarily motivated by results from simulations (see e.g. Zhang et al., 2021; Theis et al., 2024). In the subsequent sections, we present the comparisons of the six different models which support our decision.
With the simulator set to the desired configuration and priors defined in Table 1, the SBI pipeline is run with the TSNPE algorithm, which returns the posterior distribution for the free parameters. In Figure 3, we show the marginalised contour plots using ChainConsumer (Hinton, 2016). The forward simulations have the resolution parameter $N_{\mathrm{side}}=4096$. As mentioned at the beginning of this section, there are three parameters that we constrain; $\mathcal{A}$ as a scale for the input cosmology, $\mathrm{DM}_{\mathrm{host}}$ as the median and $\sigma_{\mathrm{LN}}$ as the scale of the log-normal host distribution. The value for the scale parameter $\mathcal{A}$ is $0.89^{+0.31}_{-0.29}$ at $68\%$ confidence, which is consistent with unity or the input fiducial Planck cosmology. The median and the scale of the log-normal host along with their $68\%$ confidence intervals are given as $\mathrm{DM}_{\mathrm{host}}=223^{+108}_{-95}\;\mathrm{pc\;cm}^{-3}$ and $\sigma_{\mathrm{LN}}=0.76^{+0.37}_{-0.44}$. The scale parameter essentially controls the skewness of the distribution. However, for a more intuitive comparison of the fiducial host (log-normal) to the non-fiducial host (Gaussian and Gamma) distributions, we convert $\sigma_{\mathrm{LN}}$ into the standard deviation from the definition of the variance in Equation 19. Then, we write $\sigma=263_{-113}^{+91}$. As expected, we find a strong anti-correlation between $\mathcal{A}$ and $\mathrm{DM}_{\mathrm{host}}$ since they both enter the observed DM in an additive fashion. The width of the host contribution is not degenerate with the other parameters, as it rather reduces the error on each measurement than changing the signal.
Our constraints of $\sigma_{\mathrm{LN}}$ seem to be slightly prior driven, as can be seen from the posterior hitting the prior boundaries of $\sigma_{\mathrm{LB}}$ in Figure 3. However, there is clearly some constraining power in the data on the shape of the host distribution. In general, we find excellent agreement with previous work (Macquart et al., 2020a; James et al., 2022b; Hagstotz et al., 2022b; Khrykin et al., 2024b). The most direct comparison of our work is Reischke & Hagstotz (2023a), where the scale $\sigma$ is kept constant. In our case, the mean of $\mathcal{A}$ improves by $\sim 31\%$ and the errors on $\mathrm{DM}_{\mathrm{host}}$ shrink by $\sim 22\%$ while the mean agrees at $\sim 62\%$.
@@
Additionally, we check for the goodness of fit of the model, and the result is shown in Figure 4. The histograms are the $\chi^{2}$ values from our simulations at best-fit values of the parameters, and the vertical line corresponds to the $\chi^{2}$ evaluated at the observed DM. The PTE for the fiducial model is 0.3 (corresponding to a $p$-value), which means that the observation agrees with the majority of the probability mass and is not an outlier, implying that the model is a good fit. We can thus concur that the ‘log-normal field and log-normal host’ model is an excellent fit for the DM of FRBs.
@@
With the SBI framework, all the correlations are taken into account without defining any likelihood function. Nonetheless, we can still recover the likelihood from the simulated DMs for the best-fit values. We, therefore, run our pipeline 1000 times at the best-fit value and plot the corresponding distribution of DM. In Figure 5, we show the DM likelihoods for two FRBs with varying redshift, $z=0.1178$ and $z=0.66$. The first observation is the non-Gaussian nature of the distribution, with a tendency towards higher DMs. This behaviour emerges from the host and LSS DM contribution. If the electron density of the host halo where the FRB progenitor resides is high, the signal experiences higher dispersion. That seems to be especially the case for FRB 20190520B, with an observed $\mathrm{DM}=1202\;\mathrm{pc\;cm}^{-3}$ at a relatively lower redshift of $z=0.241$ as shown in Footnote 3. The high DM is due to the local contribution of the dwarf host galaxy, identified as J160204.31-111718.5 (Ocker et al., 2022; Niu et al., 2022; Yan et al., 2024). The implication is that a Gaussian likelihood assumption in the inference can introduce biases, as it does not consider the physical effect of variable travel distances and local environments of the host. Furthermore, we observe that the mean shifts towards higher DMs in keeping with the DM-$z$ relation. We also see that the width of the distribution seems to be getting slightly smaller. This might be counter-intuitive at first, as the LSS component should increase with redshift, hence causing a broader distribution. However, the host contribution is scaled down with $1/(1+z)$ in physical coordinates. At the same time, the density of the electron density field reduces as well with increasing redshift, the DM itself, is still an additive effect. Thus, those two counter-acting effects can lead to a decreasing width of the likelihood as a function of redshift. The strength of this effect itself depends on the parameters of the model, in particular on the parameters of the host contribution. To understand this a bit better, let us consider a simple toy model. We call the variance of the host at redshift zero $\sigma^{2}_{\mathrm{host,0}}$ and label the variance of the LSS at $z=1$ as $\sigma^{2}_{\mathrm{LSS,1}}$. For linear structure growth in a matter-dominated Universe, one has $\delta_{\mathrm{e}}\propto a$. The LSS variance scales roughly linear with redshift in this case (Reischke & Hagstotz, 2023b). Therefore, the total variance is given by
@@
Finally, we show the model prediction at the global maximum posterior together with the likelihood for all data points in Figure 6. Note that the model prediction includes all components of the DM in Equation 1. Therefore, the relation between DM and redshift is not necessarily monotonous. The medians of the simulated DMs are shown in cyan dots for the redshifts of the FRBs. The colour bars represent the different percentiles. They include the observed data (black cross) within the darker shades, i.e. 25 to 75 percentiles. We would like to remark that those percentiles are, in principle, correlated due to the LSS contribution. However, this correlation is not important for the number of FRBs considered here (Reischke & Hagstotz, 2023b). It is again apparent that the long tails of the likelihood are required to explain the large scatter in DM values. We furthermore investigated the response of the inference to leaving out FRB20190520B, which has a DM > 1000 but is located at a low redshift. Our findings show that the host contribution responds with a lower median and width by roughly 10 and 20 percent, respectively. This, of course, is still fully consistent within the error bars. It shows, however, that this particular FRB increases the values of the inferred host contribution. Lastly, we recover the mean DM-$z$ relation as well by connecting the cyan dots.
@@
Figure 5: Likelihood of the DM for two FRBs at different redshifts as evaluated from the simulations. The blue line is for a lower redshift of $z=0.1178$, while the red line is for $z=0.66$. The non-Gaussian nature is obvious, along with the distinct long tail, which accounts for the high DM values originating from the disk of the host halo.
6.2 Non-fiducial model and comparison
In the following, we move from the fiducial model and compare all the possible combinations of LSS field and host distributions. With our established pipeline, this requires changing only the sampling function from ‘log-normal’ to ‘Gaussian’, for example. We also reduced the resolution parameter, $N_{\mathrm{side}}$, to 512 to decrease the computation time. This, however, produces bias or overconfidence in the TARP coverage test. We observe that this is from a combined effect of $(i)$ not including the DM contribution from smaller scales and $(ii)$ smoothening of pixels from interpolation, both of which result in lower variance of the LSS component. We elaborate more on this effect in appendix A. We note that the bias at low $N_{\mathrm{side}}$ is present even for the fiducial case and is minimised by increasing the resolution. This allows us to compare the different combinations directly and any comparative conclusion we draw is agnostic to the choice of $N_{\mathrm{side}}$.
6.2.1 Variations in the LSS component